OpenStudy (anonymous):

I need to combine terms for the following: 9t-6-6t=6 Does this equation turn into 6- 3t? Thanks, Henjin

OpenStudy (anonymous):

Oops I meant 6-3t=6

OpenStudy (sanjanap):

it actually goes woth -3t = 6-6 then., -3t = 0 finally, t=3

OpenStudy (sanjanap):

hope this helps! :)

OpenStudy (anonymous):

Yep!

OpenStudy (sanjanap):

Can you give me some medals